# Efficiently and Effectively: A Two-stage Approach to Balance Plaintext and Encrypted Text for Traffic Classification

### v2

@@ 2.3. Training Stage Two
In the second phase, the proposed EETP aims to adaptively select different branches for traffic classification, ensuring both effectiveness and efficiency. Simply, the plaintext classification model and the encrypted classification model utilize the current PLMs-based (Pre-training Language Model) techniques.
@@ 3.1. Public Dataset
First of all, The proposed EETP and comparison methods are evaluated using the public traffic dataset (Draper-Gil et al., 2016). Specifically, the ISCXVPN (Draper-Gil et al., 2016) dataset predominantly consists of network traffic data using Virtual Private Networks (VPNs), with each packet labeled by application, such as Gmail, Facebook, and others.
@@ 3.2. Collected Dataset
To further conduct an effective and efficient comparison between the SOTA models and EETP, we conduct a real-world dataset collection manually, named CHNAPP. This dataset comprises network traffic data from five widely-used Internet applications: Weibo, WeChat, QQMail, TaoBao, Youku, and QQMusic. By incorporating these popular applications, CHNAPP provides a robust basis for evaluating and contrasting the performance of EETP with leading models in real-world scenarios. This comprehensive collection ensures that the comparison is both realistic and relevant.
@@ 3.3. Evaluation Metrics
The category distribution in these tasks varies, with some being balanced and others imbalanced. To accurately assess performance, both macro F1 (Ma-F1) and micro F1 (Mi-F1) scores are utilized. The macro F1 score calculates the F1 score independently for each class and then takes the average, treating each class equally regardless of its size. This approach is particularly effective for evaluating model performance in scenarios with imbalanced class distributions, as it ensures that smaller classes are given the same consideration as larger ones. Conversely, the micro F1 score aggregates the contributions from all classes to compute a global F1 score. This method places more emphasis on the performance of larger classes, providing a measure of overall accuracy. By considering both macro F1 and micro F1 scores, we can provide a more comprehensive evaluation. Time metric means that the number of samples that the model predicts per second.
3.4. Experimental Setting & Baselines
@@
For the experimental settings, the model parameters are detailed in three aspects. Firstly, for the DPC selector, the training data comprises approximately 100,000 samples, while the testing data consists of 10,000 samples. The training process runs for 3 epochs with a learning rate of 1e-5 and a batch size of 64. Next, for the plaintext classification model, the one-hot layer dimension is set to 100, and the dimension of the 1D-convolution layer is set to (1, 10). This model is trained for 1 epoch with a learning rate of 1e-5 and a batch size of 64. Finally, for the encrypted text classification model, the BERT layer dimension is set to 768. All other parameters remain consistent with the model (Meng et al., 2022). Epoch is set to 5 , with a learning rate of 5e-6 and a batch size of 64. The Adam (Kingma and Ba, 2014) optimizer with the initial learning rate of $1$e-$5$ is utilized for training, and the Pytorch framework (Paszke et al., 2017) is adapted. The experiments are implemented on the Tesla A100-80G GPU.
@@
The baselines can be divided into three categories for comparison: (1) feature-based methods (i.e., APPS (Taylor et al., 2018), SMT (Shen et al., 2019)); (2) byte-based models (i.e., Deep Packet (Lotfollahi et al., 2020), TR-IDS (Min et al., 2018), HEDGE (Casino et al., 2019), 3D-CNN (Zhang et al., 2020)); and (3) PLMs-based methods (i.e., BERT (Devlin et al., 2019), PacRep (Meng et al., 2022), ET-BERT (Lin et al., 2022), YaTC (Zhao et al., 2023)). More details of these compared baselines can be seen in Appendix A.
@@ 3.5. Main Results
Considering the different techniques employed by feature-based methods, byte-based models, and PLMs-based methods, we focus our reproduction only on the PLMs-based methods to ensure a fair comparison of time overheads and provide time efficiency metrics for these methods. In addition, to highlight the effectiveness of PLMs-based methods, we compare their macro-F1 and micro-F1 scores with those of feature-based methods and byte-based models. The conclusions drawn from this comparison are as follows: (1) In terms of effectiveness , feature-based methods and byte-based models perform comparably, while PLMs-based methods demonstrate further improvement. Notably, our proposed EEPT model achieves SOTA results across almost all evaluation metrics, illustrating the superiority of the two-stage framework, which adaptively make a balance between the plaintext and encrypted text. (2) From a efficiency perspective, EEPT exhibits a 2 to 5 times improvement across two datasets. Specifically, when compared to the pacrep model—a BERT-based model fine-tuned directly under supervised learning—the EEPT model shows clear advantages. ET-BERT and YaTC, on the other hand, incorporate an additional pre-training phase before fine-tuning, resulting in greater overhead compared to directly fine-tuning BERT. Overall, the analysis of both effectiveness and efficiency demonstrates the superiority of the EEPT framework and underscores the necessity of analyzing both plaintext and encrypted text.
@@ 4. Conclusion
In this paper, we propose a simple but Efficient and Effective Two-Phase approach (EETP) to balance plaintext and encrypted text for traffic classification. Stage one is to determine whether the plain text information is enough to be accurately classified using the proposed DPC Selector. Stage two aims to adaptively make a classification with the result from stage one. The proposed two-phase framework provides a new insight to make an analysis on the plaintext and encrypted text in network traffic field. In summary, EETP is very effective and efficient, which outperforms the previous methods on two datasets. For future work, the design of modules in the EEPT could be considered for further research.
Appendix A Details of Baselines

### v3

@@ 3.3. Training Stage Two
In the second phase, the proposed EETS aims to adaptively select different branches for traffic classification, ensuring both effectiveness and efficiency. To further enhance the model’s efficiency, the plaintext classification model will employ a simple CNN-based deep learning approach, which allows for faster predictions. On the other hand, the encrypted classification model will utilize more complicated DL-based techniques, like attention mechanism or Pre-trained Language Model (PLM), to ensure high accuracy. Below, we provide a detailed introduction to these two modules.
3.3.1. Plaintext Classification Model
The plaintext classification model leverages CNNs to enhance the model’s efficiency, which are well-suited to quickly process and classify a large number of data. CNNs are advantageous in this context due to their efficiency in feature extraction and their ability to capture contextual relationships in the data. This allows the model to make fast and accurate predictions, making it ideal for scenarios where rapid decision-making is critical. Specifically, the model begins by tokenizing the input plaintext to obtain a sequence of strings. These words are then converted into corresponding numerical values using the one-hot encoding layer, resulting in a sequence $\{h_{1},h_{2},\dots,h_{M}\}$:
(2) $\{h_{1},h_{2},\dots,h_{M}\}={\rm{One\verb|-|hot~{}Layer}}~{}(\{x^{p}_{1},x^{p}_{2},\dots,x^{p}_{M}\})$
where $h_{m}\in\mathbb{R}^{1\times k}$, $k$ is the dimension of one-hot layer, $m$ is the index, $M$ is the total number of plaintext words.
This numerical representation enables the model to perform subsequent computations. Next, a 1D-convolution layer with a dimension of (1, $d_{c}$) is employed to capture local dependencies within the sequence and extract essential features. The advantage of using a 1D-convolution layer lies in its ability to efficiently process sequential data by sliding a filter across the input, thus detecting patterns and relationships within a localized context. The encoded hidden vectors are as follows:
(3) $\{\tilde{h_{1}},\tilde{h_{2}},\dots,\tilde{h_{M}}\}={\rm{1D\verb|-|convolution~{}Layer}}~{}(\{h_{1},h_{2},\dots,h_{M}\})$
where $\tilde{h_{m}}\in\mathbb{R}^{1\times d_{c}}$, $d_{c}$ is the dimension of 1D-convolution layer.
Following this, max-pooling is applied to reduce the dimensionality of the hidden vectors and highlight the most significant features, enhancing computational efficiency and robustness to variations in the input. Batch normalization and dropout are then utilized to normalize the activations and prevent overfitting, respectively. Finally, average pooling is performed before passing the features through a softmax layer to output the probability distribution $p\in\mathbb{R}^{l}$ over the classes, where $l$ is the total number of applications in the dataset. This distribution indicates the likelihood of the input belonging to each classes.
3.3.2. Encrypted Text Classification Model
For the encrypted classification model, a more intricate DL-based method is employed to handle the additional complexity of encrypted data. Techniques such as attention mechanism or Transformer-based models, e.g.BERT (Devlin et al., 2019), can be considered for their superior ability to manage sequential and highly structured data. These models, though computationally intensive, provide the necessary accuracy to classify encrypted traffic effectively. At the same time, after the first selecting stage, only a smaller portion of the data is towards this branch. Specifically, given the complexity of encrypted text, we utilize pre-trained language models such as BERT (Devlin et al., 2019; Lin et al., 2022) as the encoder in our encrypted classification model. The BERT is employed to encode network traffic data, resulting in rich hidden representations. The choice of BERT leverages its capability to understand and process complex traffic patterns, which is crucial for effectively handling encrypted traffic data, the equation is defined as:
(4) $\{h_{1},h_{2},\dots,h_{N}\}={\rm{BERT}}~{}(\{x^{p}_{1},x^{p}_{2},\dots,x^{e}_{N}\})$
where $h_{n}\in\mathbb{R}^{1\times d_{b}}$, $d_{b}$ is the dimension of BERT layer, $n$ is the index, $N$ is the total number of traffic words.
Subsequently, we apply an attention mechanism to model the interactions between different parts of the encoded traffic data. The attention mechanism allows the model to focus on the most relevant portions of the data, enhancing the model’s ability to capture intricate dependencies within the traffic sequences. The updated hidden representation can be expressed in the following:
(5) $\{\tilde{h_{1}},\tilde{h_{2}},\dots,\tilde{h_{N}}\}={\rm{Softmax}}~{}(\frac{QK^{T}}{\sqrt{d_{b}}})V$
where $Q=K=V=\{h_{1},h_{2},\dots,h_{N}\}$, and $\tilde{h_{n}}$ denotes the refined hidden representations post-attention.
Finally, these refined hidden representations are passed through a classifier, e.g. multilayer perceptron, which outputs the probability distribution over the possible traffic classes, as:
(6) $P(y|\tilde{H})={\rm{Softmax}}~{}({\rm{Classifier}}~{}(\{\tilde{h_{1}},\tilde{h_{2}},\dots,\tilde{h_{N}}\}))$
Table 1. Statistics of the datasets, including two public datasets and one collected real-world dataset.
Dataset Total Packets Train Valid Test
ISCXTor 614,575 450,001 82,287 82,287
ISCXVPN 492,598 443,337 24,631 24,630
CHNAPP 1,287,303 1,158,520 64,391 64,392
Table 2. Comparison results on public and collected datasets. indicates reproduced results. Effi. metric means that the number of samples that the model predicts per second. Ma-F1 and Mi-F1 display using percentages (%).
Methods ISCXVPN ISCXTor CHNAPP
Ma-F1 Mi-F1 Effi. Ma-F1 Mi-F1 Effi. Ma-F1 Mi-F1 Effi.
APPS (Taylor et al., 2018) 81.68 81.33 - - 49.68 - - - -
SMT (Shen et al., 2019) 76.31 76.00 - - 31.54 - 34.77 45.62 -
Deep Packet (Lotfollahi et al., 2020) 65.06 66.70 - - 26.81 31.28 41.07 -
TR-IDS (Min et al., 2018) 54.98 57.66 - - 20.75 - - - -
HEDGE (Casino et al., 2019) 29.67 34.00 - - 15.08 - 25.67 36.35 -
3D-CNN (Zhang et al., 2020) 54.34 57.00 - - 33.96 - 38.74 51.12 -
BERT (Devlin et al., 2019) 88.45 88.73 41.04 61.84 92.25 58.33 56.94 84.08 32.88
ET-BERT (Lin et al., 2022) 94.25 96.05 10.83 95.12 97.86 18.99 94.08 96.09 20.14
YaTC (Zhao et al., 2023) 96.67 98.04 23.68 96.30 98.15 23.07 90.44 93.50 29.41
PacRep (Meng et al., 2022) 99.32 99.13 22.75 95.02 96.38 36.76 95.10 98.75 44.46
EETS (Ours) 99.89 99.88 124.01 97.56 98.87 171.43 94.95 99.14 82.11
@@ 4.1. Public Dataset
First of all, the proposed EETS and comparison methods are evaluated using two public traffic datasets (Draper-Gil et al., 2016; Lashkari et al., 2017). Specifically, the ISCXVPN (Draper-Gil et al., 2016) dataset predominantly consists of network traffic data using Virtual Private Networks (VPNs), with each packet labeled by application, such as Gmail, Facebook, and others. The ISCXTor (Lashkari et al., 2017) dataset utilizes The Onion Router (Tor) to enhance communication privacy and includes 13 categories of application classification, like Bittorent, Netflix, Youtube and so on. Detailed statistical information on the number of packets used for training, validation, and testing is provided in Table 1.
@@ 4.2. Collected Dataset
To further conduct an effective and efficient comparison between the SOTA models and EETS, we conduct a real-world dataset collection manually, named CHNAPP. This dataset comprises network traffic data from six widely-used Internet applications: Weibo, WeChat, QQMail, TaoBao, Youku, and QQMusic. By incorporating these popular applications, CHNAPP provides a robust basis for evaluating and contrasting the performance of EETS with leading models in real-world scenarios. This comprehensive collection ensures that the comparison is both realistic and relevant. The dataset will be released in the future, and the statistical dataset information can be seen in Table 1.
@@ 4.3. Evaluation Metrics
Evaluation metrics are designed from two perspective. For effectiveness, the category distribution in these tasks varies, with some being balanced and others imbalanced. To accurately assess performance, both Macro F1 (Ma-F1) and Micro F1 (Mi-F1) scores are utilized. The Macro F1 score calculates the F1 score independently for each class and then takes the average, treating each class equally regardless of its size. This approach is particularly effective for evaluating model performance in scenarios with imbalanced class distributions, as it ensures that smaller classes are given the same consideration as larger ones. Conversely, the Micro F1 score aggregates the contributions from all classes to compute a global F1 score. This method places more emphasis on the performance of larger classes, providing a measure of overall accuracy. By considering both Macro F1 and Micro F1 scores, we can provide a more comprehensive evaluation. For the efficiency calculation, Effi. (Efficiency) metric means that the number of samples that the model predicts per second. The above evaluation metrics are the higher the value, the better.
4.4. Experimental Setting
@@
For the experimental settings, the model parameters are detailed in three aspects. Firstly, for the DPC selector, the training data comprises approximately 100,000 samples, while the testing data consists of 10,000 samples. The training process runs for 3 epochs with a learning rate of 1e-5 and a batch size of 64. Next, for the plaintext classification model, the one-hot layer dimension is set to 100, and the dimension of the 1D-convolution layer is set to (1, 10). This model is trained for 1 epoch with a learning rate of 1e-5 and a batch size of 64. Finally, for the encrypted text classification model, the BERT layer dimension is set to 768. All other parameters remain consistent with the Pacrep model (Meng et al., 2022). Epoch is set to 5 , with a learning rate of 5e-6 and a batch size of 64. The Adam (Kingma and Ba, 2014) optimizer with the initial learning rate of $1$e-$5$ is utilized for training, and the Pytorch framework (Paszke et al., 2017) is adapted. The experiments are implemented on the Tesla A100-80G GPU.
4.5. Baselines
@@
The baselines can be divided into three categories for comparison: (1) feature-based methods (i.e., APPS (Taylor et al., 2018), SMT (Shen et al., 2019)); (2) byte-based methods (i.e., Deep Packet (Lotfollahi et al., 2020), TR-IDS (Min et al., 2018), HEDGE (Casino et al., 2019), 3D-CNN (Zhang et al., 2020)); and (3) PLMs-based methods (i.e., BERT (Devlin et al., 2019), PacRep (Meng et al., 2022), ET-BERT (Lin et al., 2022), YaTC (Zhao et al., 2023)). In this paper, PLMs-based methods are main comparison points due to their competitive performance. More details of these compared baselines can be seen in Appendix A.1 because of the space limitation.
Refer to caption Refer to caption
(a) Confusion matrix on plaintext classifiable setting. (b) Confusion matrix on plaintext non-classifiable setting.
Figure 5. Confusion matrix visualization on two different settings on CHNAPP dataset. 0: QQMail, 1: QQMusic, 2: Youku, 3: TaoBao, 4: WeChat, 5: Weibo, where the proportion of QQMail data is relatively small.
Table 3. Comparison results on different settings. indicates reproduced results.
Methods ISCXVPN ISCXTor CHNAPP
Ma-F1 Mi-F1 Effi. Ma-F1 Mi-F1 Effi. Ma-F1 Mi-F1 Effi.
PacRep (Meng et al., 2022) 99.32 99.13 22.75 95.02 96.38 36.76 95.10 98.75 44.46
Plaintext Classifiable 99.98 99.97 304.41 97.16 99.97 314.40 79.85 99.89 103.40
Plaintext non-Classifiable 77.62 94.22 19.78 86.48 92.75 48.03 79.58 86.18 18.27
Overall 99.89 99.88 124.01 97.56 98.87 171.43 94.95 99.14 82.11
@@ 4.6. Main Results
Considering the different techniques employed by feature-based methods, byte-based models, and PLMs-based methods, we focus our reproduction only on PLMs-based methods to ensure a fair comparison of time overheads and provide time efficiency metric (Effi.) for these methods. In addition, to highlight the effectiveness of PLMs-based methods, we compare their Macro-F1 and Micro-F1 scores with those of feature-based methods and byte-based models. Due to the relatively poor accuracy of feature-based methods and byte-based models, we do not compare their time overheads with our approach. Instead, we focus on comparing the time efficiency of our method with the current SOTA models, i.e. PLMs-based methods, to ensure both effectiveness and efficiency.
As shown in Table 2, the conclusions drawn from this comparison are as follows: (1) In terms of effectiveness , feature-based methods and byte-based models perform comparably, while PLMs-based methods demonstrate further improvement. Notably, our proposed EETS model achieves SOTA results across almost all evaluation metrics, illustrating the superiority of the two-stage framework, which adaptively makes a balance between the plaintext and encrypted text. It is noteworthy that although EETS may obtain slightly lower performance in certain metrics, such as a 0.15% reduction in the Macro-F1 score on the CHNAPP dataset, the model achieves approximately a 2 to 4 times improvement in efficiency. This significant enhancement in time efficiency validates that the proposed method not only maintains robust effectiveness but also substantially improves the prediction time efficiency of the model.
(2) From a efficiency perspective, EETS exhibits a notable improvement across three datasets. Specifically, when compared to the PacRep model, a BERT-based model fine-tuned directly under supervised learning, the EETS model shows clear advantages (about 2 to 5 times improvement). On the other hand, ET-BERT and YaTC incorporate an additional pre-training phase before fine-tuning, resulting in greater overhead compared to directly fine-tuning BERT.
Overall, the analysis of both effectiveness and efficiency demonstrates the superiority of the EETS framework and underscores the necessity of analyzing both plaintext and encrypted text.
4.7. Results on Different Settings
To perform a fine-grained analysis of our model’s final performance (i.e., to understand why the model performs well), we conduct a detailed comparison in plaintext classifiable and non-classifiable settings. The backbone of our model is Pacrep. As shown in Table 3, our model achieves overall optimal performance across three datasets, excelling in both effectiveness and efficiency metrics.
Specifically, in the plaintext classifiable setting, our model EETS demonstrates superior prediction efficiency, outperforming baseline models by approximately 3 to 15 times while maintaining decent accuracy. This significant improvement underscores the model’s capability to handle classifiable plaintext data efficiently. In addition, in the plaintext non-classifiable setting, the model’s accuracy is relatively lower. This phenomenon is likely due to the inherent difficulty of classifying these difficult samples. Despite this, our model still achieves almost SOTA results in terms of overall performance. For further analysis of the Macro-F1 scores, refer to Section 5.1 Confusion Matrix Visualization.
Table 4. Results of DPC selector analysis on different datasets. Num.: the number of the samples.
Methods ISCXTor ISCXVPN CHNAPP
Ma-F1 Mi-F1 Num. Ma-F1 Mi-F1 Num. Ma-F1 Mi-F1 Num.
DPC Selector (ours) 97.56 97.56 82,287 89.65 99.61 24,630 92.45 98.60 64,392
Refer to caption Refer to caption
(a) Testing sample size analysis on setting one. (b) Testing sample size analysis on setting two.
Figure 6. Testing sample size analysis on two different settings.
5. Analysis
5.1. Confusion Matrix Visualization
In Table 3, we have made a new discovery. Horizontally examining the three datasets, particularly CHNAPP, we observe that the Macro-F1 scores under both settings are relatively low. However, the final Macro-F1 score reaches an impressive 94.95%. To analyze this phenomenon, we conduct a detailed statistical analysis of the results for different classes within CHNAPP and visualize the confusion matrix in Figure 5. Figures 5 (a) and (b) represent the plaintext classifiable and non-classifiable settings, respectively. From Figure 5 (a), it is evident that most samples are accurately classified. Specifically, the accuracy for the last four classes is nearly 99%, although there is some misclassification in the QQmail class, which leads to a lower Macro-F1 score (79.85%), while the Micro-F1 score reaches an outstanding 99.89%. Additionally, in Figure 5 (b), we observe: (1) The classification difficulty in setting two is higher compared to the plaintext classifiable setting. Specificially, there exists lower performance on the Youku class samples, while other categories still achieve good levels of accuracy. (2) The QQmail class samples are all correctly classified, further proving that the introduction of encrypted text effectively helps the model in classification. This makes up for the misclassification issue of the QQmail class samples observed in setting one.
All in all, the proposed model’s performance metrics on the CHNAPP dataset are decent, with a Macro-F1 of 94.95% and a Micro-F1 of 99.14%. This observation holds true for other datasets as well, indicating consistent performance improvements across different scenarios.
5.2. DPC Selector Analysis
The proposed method is a two-stage approach where the outcome of the first stage serves as the input for the second stage. Consequently, the performance of the model in the first stage significantly impacts the results of the second stage. To further analyze the model’s performance and demonstrate the effectiveness of the two-stage approach, this section presents the DPC selector analysis. Specifically, we examine the performance of the DPC selector in the first stage. As shown in the Table 4, the Micro-F1 scores exceed 97% across all three datasets, indicating that almost all data can be correctly selected, thereby ensuring robust performance in the second stage.
Regarding the Macro-F1 metric, the score for the ISCXTor dataset is slightly lower at 89.65%, likely due to data imbalance. However, the Micro-F1 score for ISCXTor reaches 99.61%, making the impact on the final classification results negligible. Additionally, the Macro-F1 scores remain high for the other two datasets. All in all, this experimental analysis confirms the effectiveness of the proposed DPC selector, therefore guaranteeing the accurate traffic classification in the second stage.
5.3. Testing Sample Size Analysis
To further validate the performance of the DPC selector and the reliability of plaintext classification, we conduct a sample size analysis under both plaintext classification and non-classification settings. As shown in Figure 6 (a), in setting one, the ratio of samples predicted to be plaintext-classifiable closely matches the actual samples, with both ratios covering around 50%. Similarly, in setting two, the ratio is maintained between 45% and 55%, with a slightly larger variance observed in the ISCXVPN dataset. These results further demonstrate the effectiveness of the DPC selector.
In a cross-comparison, as illustrated in Figures 6 (a) and (b), the number of samples predicted to be plaintext-classifiable accounts for a large proportion. Specifically, in the ISCXVPN dataset, the results reach nearly 100 times (24,439/191), while in the other two datasets, the results range from 7 to 30 times. This strongly indicates that relying on plaintext information can address the majority of traffic classification scenarios. And encrypted data can also aid in improving classification accuracy alongside plaintext. Moreover, the simplicity and readability of plaintext further enhance the model’s prediction efficiency, providing that the motivation behind this paper is both effective and necessary.
@@ 6. Conclusion
In this paper, we propose a simple but Efficient and Effective Two-Stage approach (EETS) to balance plaintext and encrypted text for traffic classification. Moreover, we make the first attempt to analyze the impact of plaintext and encrypted text on model performance and time efficiency, which provides a new insight to make an analysis on the plaintext and encrypted text in network traffic field. For experiments, EETS is very effective and efficient, which outperforms the previous methods on two public datasets and one real-world dataset collected by ourselves. The analyses also demonstrate the necessity and interpretability. For future work, the design of modules in the EETS could be considered for further research.
Appendix A Appendix
A.1. Details of Baselines